Mix a solution that is 30% alcohol with a solution that is 80% alcohol to make 800 mL of a solution that is 60 % alcohol. How much of each solution should you use?
Answer by checkley77(12844)

.80x+.30(800-x)=.60*800
.80x+240-.30x=480
.50x=480-240
.50x=240
x=240/.5
x=480 ml of 80% solution is used.
800-480=320 ml of 30% solution is used.
Proof:
.80*480+.30*320=.60*800
384+96=480
480=480